Home‎ > ‎Crew 144‎ > ‎

1203-CommandersReport

Commander's Report

12/03/2014

Gregory Leonard

All crew reported very good sleeping last night, perhaps due to a very calming pre lights-out mindfulness meditation session led by HSO Dr. Susan Jewell. No doubt, such activities will prove essential for crew well-being on long duration missions in confined and isolated spaces, such as at FMARS in the high Arctic, and eventually on Mars.

 

There was also apparently a few minutes of light rain last night, although it was insufficient to soak the terrain around the Hab and thwart our scheduled EVA. Therefore after a hearty group breakfast and daily plan briefing, the team supported the preparations for EVA-2. Crew Biologist Yusuke Murakami presented his EVA plan and field objectives and directed his EVA team members Engineer Victor Luo and HSO/Journalist Susan Jewell with their specific EVA tasks. Yusuke’s investigation and outreach project is an observational analog survey characterizing geophysical zones where extra-terrestrial life could exist (see Science Report).

 

The EVA team was out onto the surface at 1002 and completed their field plans using EVA vehicles, and also by foot. Upon return to the Hab, engineering EVA duties were completed. Post-EVA debriefing involving all mission crew centered on the following issues: 1) ventilation fan on Pack 3 failed after two hours, apparently losing battery power, this despite having left under full charge. We will troubleshoot the problem, and use Pack 4 in its place; Pack 3 may still be used for shorter duration EVAs, close to the Hab. 2) some condensation occurred in all helmets to varying degrees, apparently due to the warm day and perhaps some increased humidity resultant from the recent influx of a regional moist air mass. We replaced a silica desiccant pack in one of the helmets which may reduce some condensate build up. 3) HSO Jewell receive a small bump to the back of her head when she had a minor stumble (non-fall), and her helmet struck her. Post-EVA inspection showed that the helmet contained a tiny exposed bolt that likely caused the bump to her head. The tiny bolt fastens the rear clip on the helmet collar. The bolt has now been covered with a small plastic cap and some foam, secured by duct tape. Other helmets were inspected and they all had some type of foam protection.

 

GreenHab Officer Heidi Beemer activated the hydroponics system which was developed and built by a previous crew. She reported leakage in the system which requires further identification and repairs (see GreenHab Report).

 

Crew Scientist Juho Vehviläinen assisted me in rehearsals and EVA field operational sequencing for my upcoming EVA work. The investigation involves complex sensor-data logger- software interfacing and calibration, and field measurement protocols. Juho has worked with very similar spectroradiometer instruments, and has considerable field experience in remote and rugged terrains as well; he was therefore quite invaluable to me today.

 

Crew Geologist Toni de Morais Teles continued photo-documenting the near-Hab landscape, and preparing his equipment for his forthcoming EVA work. In addition he singlehandedly cleaned the kitchen/mess/ working space in the upper Hab deck, which is now spotless. We all greatly recognize the importance of regular cleaning of the Hab; it promotes good and essential Hab hygiene, which is indeed important since the Hab is for this time, our precious and only home.

 

Our HSO continues to monitor two crew members for illness (see HSO report).

 

The spirit of this crew is something remarkable. All crew are respectful, generous, and truly contributory towards each other and our mission. It is gratifying to observe this team as we continue to form stronger bonds together, forged through our common mission vision, hard work, and plenty of spontaneous discussions and laughter. Goodnight from our Mars.